# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ Appendix D Technical Lemmas
$\mathbf{F}$ $=(\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}(\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$
$+(\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}+\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}(\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}).$
Next, by Lemma C.1, we have
$\begin{split}\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}&=\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{i}}^{\top}\mathbf{X}_{\mathcal{U}_{i}})^{-1}+\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\\ &=\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{U}^{(i)}(\bm{\Lambda}^{(i)})^{-1/2}\tilde{\mathbf{W}}^{(i)}+\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle},\end{split}$
where $\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ and $\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ contain the rows in $\mathbf{E}^{(i)}$ and $\mathbf{R}^{(i)}$ corresponding to entities $\mathcal{U}_{i}\cap\mathcal{U}_{j}$, respectively. A similar expansion holds for $\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$. We therefore have
$\mathbf{F}$ $=\underbrace{(\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}(\hat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}_{\mathbf{F}_{1}}$
$+\underbrace{\tilde{\mathbf{W}}^{(i)\top}(\bm{\Lambda}^{(i)})^{-1/2}\mathbf{U}^{(i)\top}\mathbf{E}^{(i)\top}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}}_{\mathbf{F}_{2}}+\underbrace{\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{E}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{U}^{(j)}(\bm{\Lambda}^{(j)})^{-1/2}\tilde{\mathbf{W}}^{(j)}}_{\mathbf{F}_{3}}$
@@
For $\mathbf{F}_{1}$, by Lemma C.1 we have
$\|\mathbf{F}_{1}\|$ $\leq n_{i,j}^{1/2}\|\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\cdot n_{i,j}^{1/2}\|\hat{\mathbf{X}}^{(j)}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}$
@@
with high probability. For $\mathbf{F}_{2}$, by Lemma D.10 we have
@@
with high probability. For $\mathbf{F}_{4}$, once again by Lemma D.10 we have
@@
with high probability. Substituting the above bound for $\|\mathbf{F}\|$ into Eq. (D.28) yields the stated claim.
D.6 Technical lemmas for Lemma C.2
Lemma D.10.
@@
From Lemma C.1 we have
@@
with high probability. Furthermore, following the same derivations as that for $\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|$ in the proof of Lemma D.1, we have
@@
$\lambda_{i,\min}\geq\lambda_{\min}(\mathbf{U}_{\mathcal{U}_{i}}^{\top}\mathbf{U}_{\mathcal{U}_{i}})\cdot\lambda_{d}(\bm{\Lambda})\gtrsim\frac{n_{i}}{N}\lambda_{\min},$ (D.29)
@@
$\lambda_{i,\max}=\|\mathbf{P}^{(i)}\|\leq\|\mathbf{U}_{\mathcal{U}_{i}}\|^{2}\cdot\|\bm{\Lambda}\|\lesssim\frac{dn_{i}}{N}\lambda_{\max}$ (D.30)
@@
D.7 Proof of Lemma C.3
We proceed by induction on $L$. The case $L=1$ follows from Lemma C.2. Now for any $L\geq 2$, define
@@
and suppose the stated bound holds for $L-1$, i.e.
$\|\mathbf{T}^{(L-1)}-\mathbf{I}\|$ $\lesssim\sum_{\ell=1}^{L-1}\alpha_{{i_{\ell-1}},{i_{\ell}}}$ (D.31)
@@
and hence, by combining Eq. (D.31) and Lemma C.2 (for $\mathbf{W}^{(i_{L-1})\top}\mathbf{W}^{(i_{L-1},i_{L})}\mathbf{W}^{(i_{L})}-\mathbf{I}$), we obtain
@@
D.8 Extension to symmetric indefinite matrices
We first state an analogue of Lemma C.1 for symmetric but possibly indefinite matrices.
Lemma D.12.
@@ D.8 Extension to symmetric indefinite matrices
$\mathbf{P}^{(i)}=\mathbf{U}^{(i)}\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top},\quad\mathbf{A}^{(i)}=\hat{\mathbf{U}}^{(i)}\hat{\bm{\Lambda}}^{(i)}\hat{\mathbf{U}}^{(i)\top}+\hat{\mathbf{U}}^{(i)}_{\perp}\hat{\bm{\Lambda}}^{(i)}_{\perp}\hat{\mathbf{U}}^{(i)\top}_{\perp}.$
Let $d_{+}$ and $d_{-}$ denote the number of positive and negative eigenvalues of $\mathbf{P}^{(i)}$, respectively, and denote $d=d_{+}+d_{-}$. Let $\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}|\hat{\bm{\Lambda}}^{(i)}|^{1/2}$ and $\mathbf{X}^{(i)}=\mathbf{U}^{(i)}|\bm{\Lambda}^{(i)}|^{1/2}$. Suppose that
@@
$\hat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}=\mathbf{E}^{(i)}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\mathbf{I}_{d_{+},d_{-}}+\mathbf{R}^{(i)},$ (D.32)
@@
$\|\hat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}\|_{2\to\infty}\lesssim\|\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|_{2\to\infty}\cdot\|(\bm{\Lambda}^{(i)})^{-1/2}\|+\|\mathbf{R}^{(i)}\|_{2\to\infty}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\sigma_{i,\min}^{1/2}}=\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$
@@
The proof of Lemma D.12 follows the same argument as that for Lemma C.1 and is thus omitted. The main difference between the statements of these two results is that Lemma C.1 bounds $\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}$ while Lemma D.12 bounds $\hat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}$. If $\mathbf{P}$ is positive semidefinite then $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\tilde{\mathbf{W}}^{(i)}$ for some orthogonal matrix $\tilde{\mathbf{W}}^{(i)}$ and thus we can combine both $\mathring{\mathbf{W}}^{(i)}$ and $\tilde{\mathbf{W}}^{(i)}$ into a single orthogonal transformation $\mathbf{W}^{(i)}$; see the argument in Section D.2. In contrast, if $\mathbf{P}$ is indefinite then $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\mathbf{Q}^{(i)}$ for some indefinite orthogonal $\mathbf{Q}^{(i)}$. As indefinite orthogonal matrices behave somewhat differently from orthogonal matrices, it is simpler to consider $\mathring{\mathbf{W}}^{(i)}$ and $\mathbf{Q}^{(i)}$ separately.
Lemma D.13.
Consider the setting of Theorem 4 for just two overlapping submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$. Let $\mathring{\mathbf{W}}^{(i,j)}=(\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)\top}\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{-1}\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)\top}\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}=(\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{\dagger}\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ be the least square alignment between $\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. Here $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse of a matrix. Also notice $\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=(\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{\dagger}\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ is the corresponding alignment between $\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. We then have
@@
For ease of notation, we will let $\hat{\mathbf{Y}}:=\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}\mathring{\mathbf{W}}^{(i)}$ and $\hat{\mathbf{Z}}:=\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}\mathring{\mathbf{W}}^{(j)}$, and define $\mathbf{Y}:=\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\mathbf{Z}:=\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. We then have
$\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=\hat{\mathbf{Y}}^{\dagger}\hat{\mathbf{Z}}-\mathbf{Y}^{\dagger}\mathbf{Z}=(\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})(\hat{\mathbf{Z}}-\mathbf{Z})+\mathbf{Y}^{\dagger}(\hat{\mathbf{Z}}-\mathbf{Z})+(\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})\mathbf{Z},$ (D.33)
@@
Now under the assumption $\frac{n_{i,j}^{1/2}\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}\ll\theta_{i,j}^{1/2}$, we have $\|\hat{\mathbf{Y}}-\mathbf{Y}\|\ll\sigma_{d}(\mathbf{Y})$ with high probability, and hence both $\mathbf{Y}^{\top}\mathbf{Y}$ and $\hat{\mathbf{Y}}^{\top}\hat{\mathbf{Y}}$ are invertible with high probability. We thus have $\hat{\mathbf{Y}}^{\dagger}\hat{\mathbf{Y}}=\mathbf{Y}^{\dagger}\mathbf{Y}=\mathbf{I}$ and
$\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}=\hat{\mathbf{Y}}^{\dagger}(\mathbf{I}-\mathbf{Y}(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\mathbf{Y}^{\top})-\hat{\mathbf{Y}}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}.$
@@
$(\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})\mathbf{Z}=-\hat{\mathbf{Y}}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}=(\mathbf{Y}^{\dagger}-\hat{\mathbf{Y}}^{\dagger})(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}-\mathbf{Y}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}.$ (D.34)
Combining Eq. (D.33) and Eq. (D.34) we obtain
$\mathring{\mathbf{W}}^{(i)\top}\mathbf{T}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=(\hat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})(\hat{\mathbf{Z}}-\mathbf{Z})+\mathbf{Y}^{\dagger}(\hat{\mathbf{Z}}-\mathbf{Z})+(\mathbf{Y}^{\dagger}-\hat{\mathbf{Y}}^{\dagger})(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}-\mathbf{Y}^{\dagger}(\hat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z},$

### v4

@@ Appendix E Technical Lemmas
$\mathbf{F}$ $=(\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}(\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})$
$+(\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}+\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}(\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}).$
Next, by Lemma D.1, we have
[…]
where $\mathbf{E}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ and $\mathbf{R}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}$ contain the rows in $\mathbf{E}^{(i)}$ and $\mathbf{R}^{(i)}$ corresponding to entities $\mathcal{U}_{i}\cap\mathcal{U}_{j}$, respectively. A similar expansion holds for $\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$. We therefore have
$\mathbf{F}$ $=\underbrace{(\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})^{\top}(\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}})}_{\mathbf{F}_{1}}$
$+\underbrace{\widetilde{\mathbf{W}}^{(i)\top}(\bm{\Lambda}^{(i)})^{-1/2}\mathbf{U}^{(i)\top}\mathbf{E}^{(i)\top}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}}_{\mathbf{F}_{2}}+\underbrace{\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{E}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{U}^{(j)}(\bm{\Lambda}^{(j)})^{-1/2}\widetilde{\mathbf{W}}^{(j)}}_{\mathbf{F}_{3}}$
@@
For $\mathbf{F}_{1}$, by Lemma D.1 we have
$\|\mathbf{F}_{1}\|$ $\leq n_{i,j}^{1/2}\|\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\cdot n_{i,j}^{1/2}\|\widehat{\mathbf{X}}^{(j)}\mathbf{W}^{(j)}-\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}$
@@
with high probability. For $\mathbf{F}_{2}$, by Lemma E.10 we have
@@
with high probability. For $\mathbf{F}_{4}$, once again by Lemma E.10 we have
@@
with high probability. Substituting the above bound for $\|\mathbf{F}\|$ into Eq. (E.28) yields the stated claim.
E.6 Technical lemmas for Lemma D.2
Lemma E.10.
@@
From Lemma D.1 we have
@@
with high probability. Furthermore, following the same derivations as that for $\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|$ in the proof of Lemma E.1, we have
@@
$\lambda_{i,\min}\geq\lambda_{\min}(\mathbf{U}_{\mathcal{U}_{i}}^{\top}\mathbf{U}_{\mathcal{U}_{i}})\cdot\lambda_{d}(\bm{\Lambda})\gtrsim\frac{n_{i}}{N}\lambda_{\min},$ (E.29)
@@
$\lambda_{i,\max}=\|\mathbf{P}^{(i)}\|\leq\|\mathbf{U}_{\mathcal{U}_{i}}\|^{2}\cdot\|\bm{\Lambda}\|\lesssim\frac{dn_{i}}{N}\lambda_{\max}$ (E.30)
@@
E.7 Proof of Lemma D.3
We proceed by induction on $L$. The case $L=1$ follows from Lemma D.2. Now for any $L\geq 2$, define
@@
and suppose the stated bound holds for L1L-1, i.e.
$\|\mathbf{T}^{(L-1)}-\mathbf{I}\|$ $\lesssim\sum_{\ell=1}^{L-1}\alpha_{{i_{\ell-1}},{i_{\ell}}}$ (E.31)
@@
and hence, by combining Eq. (E.31) and Lemma D.2 (for $\mathbf{W}^{(i_{L-1})\top}\mathbf{W}^{(i_{L-1},i_{L})}\mathbf{W}^{(i_{L})}-\mathbf{I}$), we obtain
@@
E.8 Extension to symmetric indefinite matrices
We first state an analogue of Lemma D.1 for symmetric but possibly indefinite matrices.
Lemma E.12.
@@ E.8 Extension to symmetric indefinite matrices
$\mathbf{P}^{(i)}=\mathbf{U}^{(i)}\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top},\quad\mathbf{A}^{(i)}=\widehat{\mathbf{U}}^{(i)}\widehat{\bm{\Lambda}}^{(i)}\widehat{\mathbf{U}}^{(i)\top}+\widehat{\mathbf{U}}^{(i)}_{\perp}\widehat{\bm{\Lambda}}^{(i)}_{\perp}\widehat{\mathbf{U}}^{(i)\top}_{\perp}.$
Let $d_{+}$ and $d_{-}$ denote the number of positive and negative eigenvalues of $\mathbf{P}^{(i)}$, respectively, and denote $d=d_{+}+d_{-}$. Let $\widehat{\mathbf{X}}^{(i)}=\widehat{\mathbf{U}}^{(i)}|\widehat{\bm{\Lambda}}^{(i)}|^{1/2}$ and $\mathbf{X}^{(i)}=\mathbf{U}^{(i)}|\bm{\Lambda}^{(i)}|^{1/2}$. Suppose that
@@
$\widehat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}=\mathbf{E}^{(i)}\mathbf{X}^{(i)}(\mathbf{X}^{(i)\top}\mathbf{X}^{(i)})^{-1}\mathbf{I}_{d_{+},d_{-}}+\mathbf{R}^{(i)},$ (E.32)
@@
$\|\widehat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}\|_{2\to\infty}\lesssim\|\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|_{2\to\infty}\cdot\|(\bm{\Lambda}^{(i)})^{-1/2}\|+\|\mathbf{R}^{(i)}\|_{2\to\infty}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\sigma_{i,\min}^{1/2}}=\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$
@@
The proof of Lemma E.12 follows the same argument as that for Lemma D.1 and is thus omitted. The main difference between the statements of these two results is that Lemma D.1 bounds $\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}$ while Lemma E.12 bounds $\widehat{\mathbf{X}}^{(i)}\mathring{\mathbf{W}}^{(i)}-\mathbf{X}^{(i)}$. If $\mathbf{P}$ is positive semidefinite then $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\widetilde{\mathbf{W}}^{(i)}$ for some orthogonal matrix $\widetilde{\mathbf{W}}^{(i)}$ and thus we can combine both $\mathring{\mathbf{W}}^{(i)}$ and $\widetilde{\mathbf{W}}^{(i)}$ into a single orthogonal transformation $\mathbf{W}^{(i)}$; see the argument in Section E.2. In contrast, if $\mathbf{P}$ is indefinite then $\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{X}^{(i)}\mathbf{Q}^{(i)}$ for some indefinite orthogonal $\mathbf{Q}^{(i)}$. As indefinite orthogonal matrices behave somewhat differently from orthogonal matrices, it is simpler to consider $\mathring{\mathbf{W}}^{(i)}$ and $\mathbf{Q}^{(i)}$ separately.
Lemma E.13.
Consider the setting of Theorem C.1 for just two overlapping submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$. Let $\mathring{\mathbf{W}}^{(i,j)}=(\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)\top}\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{-1}\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)\top}\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}=(\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{\dagger}\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ be the least square alignment between $\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. Here $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse of a matrix. Also notice $\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=(\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)})^{\dagger}\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ is the corresponding alignment between $\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. We then have
@@
For ease of notation, we will let $\widehat{\mathbf{Y}}:=\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}\mathring{\mathbf{W}}^{(i)}$ and $\widehat{\mathbf{Z}}:=\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}\mathring{\mathbf{W}}^{(j)}$, and define $\mathbf{Y}:=\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\mathbf{Z}:=\mathbf{X}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$. We then have
$\mathring{\mathbf{W}}^{(i)\top}\mathring{\mathbf{W}}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=\widehat{\mathbf{Y}}^{\dagger}\widehat{\mathbf{Z}}-\mathbf{Y}^{\dagger}\mathbf{Z}=(\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})(\widehat{\mathbf{Z}}-\mathbf{Z})+\mathbf{Y}^{\dagger}(\widehat{\mathbf{Z}}-\mathbf{Z})+(\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})\mathbf{Z},$ (E.33)
@@
Now under the assumption $\frac{n_{i,j}^{1/2}\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}\ll\theta_{i,j}^{1/2}$, we have $\|\widehat{\mathbf{Y}}-\mathbf{Y}\|\ll\sigma_{d}(\mathbf{Y})$ with high probability, and hence both $\mathbf{Y}^{\top}\mathbf{Y}$ and $\widehat{\mathbf{Y}}^{\top}\widehat{\mathbf{Y}}$ are invertible with high probability. We thus have $\widehat{\mathbf{Y}}^{\dagger}\widehat{\mathbf{Y}}=\mathbf{Y}^{\dagger}\mathbf{Y}=\mathbf{I}$ and
$\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger}=\widehat{\mathbf{Y}}^{\dagger}(\mathbf{I}-\mathbf{Y}(\mathbf{Y}^{\top}\mathbf{Y})^{-1}\mathbf{Y}^{\top})-\widehat{\mathbf{Y}}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}.$
@@
$(\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})\mathbf{Z}=-\widehat{\mathbf{Y}}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}=(\mathbf{Y}^{\dagger}-\widehat{\mathbf{Y}}^{\dagger})(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}-\mathbf{Y}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}.$ (E.34)
Combining Eq. (E.33) and Eq. (E.34) we obtain
$\mathring{\mathbf{W}}^{(i)\top}\mathbf{T}^{(i,j)}\mathring{\mathbf{W}}^{(j)}-\mathbf{Q}^{(i)}(\mathbf{Q}^{(j)})^{-1}=(\widehat{\mathbf{Y}}^{\dagger}-\mathbf{Y}^{\dagger})(\widehat{\mathbf{Z}}-\mathbf{Z})+\mathbf{Y}^{\dagger}(\widehat{\mathbf{Z}}-\mathbf{Z})+(\mathbf{Y}^{\dagger}-\widehat{\mathbf{Y}}^{\dagger})(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z}-\mathbf{Y}^{\dagger}(\widehat{\mathbf{Y}}-\mathbf{Y})\mathbf{Y}^{\dagger}\mathbf{Z},$